Be sure to answer all parts. Liquid hydrogen peroxide, an oxidizing agent in many rocket fuel mixtures, releases oxygen gas on decomposition: 2 H202(1) >2 H,0(l) + 02(g) How much heat is released when 470. kg of H20, decomposes? 1.36 x 10kJ (Enter your answer in scientific notation.) You are out of attempts for this question.

Can anyone show me the correct answer?
Answer #1

Answer- M

Moles of hydrogen peroxide = 470000/34 = 13823.5 mol

Total Heat released = -196.1×13823.5/2

= -1.36×106 kJ